1. What is the present value of $1,000 that will be paid to you in two years if the interest rate were 9 percent?
Answer #1

Present value will be 841.68 $

Explanation -

Amount to be paid in 2 years = 1000 $

Interest Rate = 9%

Present value = Amount to be paid in 2 year / (1 + Interest Rate)^n

= 1000 / (1 + 9%)^2

= 1000 / (1.09)^2

= 1000 / 1.1881

= 841.68 $